RC meeting and Run Plan, Tuesday February 7th
From SBS wiki
Return to the RC Meeting Calendar
Contents
RC Report
- https://logbooks.jlab.org/entry/4126183
- RC: Rachel Montgomery, rmontgom@jlab.org
Summary
- After RC meeting yesterday we continued waiting for ops accelerator studies to finish and beam recovery.
- Started beam checkout procedure at around 1am. We were unable to re-establish beam into the Hall. Beam profile was too wide and OPS were having trouble controlling/reproducing Hall A optics.
- Today the beam is down to investigate.
- They started trying harp scans at around 1:30pm today, but then the RF separators tripped. They are working on it.
Current Status
- Waiting on to do full beam checkout when machine is ready.
Run Plan
- When ops are ready to send us beam. Do full beam checkout and then continue production running on polHe3, hopefully 45uA.
Data Collection and Polarisaton Status
- 24.0C of uncorrected charge for this kinematic (26.4C of corrected charge)
- Last NMR 52.2% (no beam)
Other Topics/Status Updates
- SBS GEM DAQ de-bugging is underway. The plan is to test put SBS GEMs in DAQ during day shifts in future.
- Anything else?
Shift Schedule
- Shift Schedule Page
- Only 9 shift spots open between owl 17th Feb and owl March 20th. All open slots are TO - please encourage someone to sign up - especially someone who only has 6 shifts!
Requested Accesses to the Hall
- Time of next access opportunity currently not clear
- Jack would like to collect some cables
- GEM team need like a 3-4hour access for the SBS GEMs DAQ troubleshooting
Special Requests
- ?
Notices for Shift Crew
- Check the How To page
- Pay attention to the beam restoration procedures in the How To
- there is a procedure if beam has been away for 30minutes, and another procedure if machine has been down for tuning or something which may change beam profile. Communicate with MCC to figure out why beam has been down and to figure out which kind of restoration you should do.
- SL:
- production runs (1-1.5hr) and replays (50k replay for each run, ~two runs also with 200k replays per shift if possible)
- TOs:
- NMR every 4 hours
- Note the instructions for performing an NMR if there has been any changes to magnetic fields (https://logbooks.jlab.org/entry/4124321)
- Contact RC any time for any questions